# Extending Stable and Popular Matching Algorithms from Bipartite to Arbitrary Instances

### v1

@@ 3.3 Popular Maximum Weight Fractional-Matching
$\begin{array}[]{rr@{}ll}\text{maximize}&\displaystyle\sum\limits_{e\in E}\omega_{e}x_{e}&&\\ \text{subject to}&\displaystyle\sum\limits_{e\in E(v)}x_{e}&\leq 1,&v\in V\\ &x_{e}&\geq 0&e\in E\\ \end{array}$
@@
$\begin{array}[]{rr@{}ll}\text{minimize}&\displaystyle\sum\limits_{v\in V}y_{v}&&\\ \text{subject to}&y_{u}+y_{v}&\geq\omega_{e}&e\in E\\ &y_{v}&\geq 0&v\in V\\ \end{array}$
@@
Let $E^{\prime}:=\{e\in E\mid y_{u}^{*}+y_{v}^{*}=\omega(e)\}$ and $\mathcal{C}:=\{v\in V\mid y_{v}^{*}>0\}$. Then, we have that the maximum weight fractional-matchings (so each possible candidate) are exactly the ones that have positive value only on edges from $E^{\prime}$ and saturate every $v\in\mathcal{C}$. Hence, we can restrict the set of edges of $G$ to $E^{\prime}$, so from now on let $E:=E^{\prime}$.
This way, we reduced the problem to finding a popular critical fractional-matching , which to recall is a fractional-matching $M$ that saturates every $v\in\mathcal{C}$ and $\Delta(M,N)\geq 0$ for any fractional-matching $N$ that saturates each $v\in\mathcal{C}$.
Therefore, now it is enough to solve Pop-Crit-SRI.
We describe the fine-tuned first step of the meta-algorithm. Let $|\mathcal{C}|=s$. We keep each edge of $E$. These will be called the middle copies. If $e=(v_{i},v_{j})$ such that $v_{i},v_{j}\notin\mathcal{C}$, we do not create additional parallel edges. If $v_{i}\in\mathcal{C},v_{j}\notin\mathcal{C}$, we add copies $e_{1},\dots,e_{s}$. These are considered worst copies for $v_{i}$ in the opposite order (i.e. $e_{s}$ is the absolute worst and $e_{1}$ is best among them) and best copies for $v_{j}$ in the opposite order (i.e. $e_{s}$ is the absolute best and $e_{1}$ is the worst among them). If $v_{i}\notin\mathcal{C},v_{j}\in\mathcal{C}$, we add copies $e_{-1},\dots,e_{-s}$. These are considered worst copies for $v_{j}$ in the opposite order (i.e. $e_{-s}$ is the absolute worst and $e_{-1}$ is best among them) and best copies for $v_{j}$ in the opposite order (i.e. $e_{s}$ is the absolute best and $e_{1}$ is the worst among them). If $v_{i},v_{j}\in\mathcal{C}$, then we do both of the above.
Then, the preferences are created such that the best copies are best, then the middle copies and then the worst copies are worst. With respect to that, each vertex ranks the edges of the same type according to their original order. For a vertex $v$, we use the notation $B_{j}^{v}(e)$ to denote the $j$-th best copy of $e$ among the best copies, and $W_{j}^{v}(e)$ to denote the $j$-th worst copy of $e$ among the worst copies (both can be either $e_{j}$ or $e_{-j}$). We also use the notation $\mathrm{lev}_{v}(B_{j}^{v}(e))=j$, $\mathrm{lev}_{v}(e)=0$ and $\mathrm{lev}_{v}(W_{j}^{v}(e))=-j$, which is called the level of the given copy of $e$ for $v$.
@@
Pop-Crit-SRI is solvable in polynomial time. Furthermore, the algorithm outputs an extra-popular matching among the critical matchings.
@@
Then, we have that $\Delta(M,N,\phi)=\sum\limits_{l=1}^{k}w_{k}\cdot\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))$. Therefore, it is enough to show that for each walk $P^{l}$, $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))\geq 0$.
@@
    If $e_{i}$ has a level $p$ copy for $v_{i}$ in $M^{\prime}$ with positive value, then if $e_{i+1}$ has a level $q$ copy for $v_{i+1}$ with positive value in $M^{\prime}$, then $q\leq p+1$.
@@
    If $e_{0}=\emptyset$, then each copy of $e_{1}$ with positive value in $M^{\prime}$ has level at most 00 for $v_{1}$ and if there is one with level exactly 00, then $\mathrm{vote}_{u_{1}}(e_{1},f_{1})=+1$.
@@
    If $e_{0}\neq\emptyset$ and $e_{0}\neq e_{j_{l}}$, then each copy of $e_{0}$ with positive value in $M^{\prime}$ has level at least 00 for $v_{0}$.
@@
    If $e_{j_{l}}=\emptyset$, then each copy of $e_{j_{l}-1}$ with positive value in $M^{\prime}$ has level at least 00 for $v_{j_{l}-1}$ and if there is one with level exactly 00, then $\mathrm{vote}_{v_{j_{l}-1}}(e_{j_{l}-1},f_{j_{l}})=+1$.
@@
    If $e_{j_{l}}\neq\emptyset$ and $e_{j_{l}}\neq e_{0}$, then each copy of $e_{j_{l}}$ with positive value in $M^{\prime}$ has level at least 00 for $v_{j_{l}}$.
@@
1. Suppose that $q>p+1$. Then, the copy of $f_{i+1}$ that has level $p+1$ for $v_{i}$ blocks $M^{\prime}$, because this copy is level $-(p+1)>-q$ for $u_{i+1}$, so both of them strictly prefer it to some copies of $e_{i}$ and $e_{i+1}$ with positive value in $M^{\prime}$, contradiction.
@@
2. In this case, $u_{i+1}$ has a level $-(p+1)$ copy of $f_{i+1}$ with positive value in $M^{\prime}$. Hence, if either $v_{i}$ or $u_{i+1}$ prefer $f_{i+1}$ to $e_{i}$ or $e_{i+1}$ respectively, then either the level $p$ or the level $p+1$ copy of $f_{i+1}$ for $v_{i}$ blocks $M^{\prime}$, contradiction.
@@
3. Suppose both $v_{i}$ and $u_{i+1}$ prefer $f_{i+1}$ to $e_{i}$ and $e_{i+1}$ respectively. Then, by the assumption of case 3, the level $p$ copy of $f_{i+1}$ for $v_{i}$ blocks $M^{\prime}$, contradiction.
@@
4. Suppose $e_{0}=\emptyset$. Then, $\phi(f_{1},\emptyset)>0$, so $v_{0}$ is unsaturated in $M^{\prime}$ by the third axiom on sensible pairings. By the fact that the level 00 copy (which must exist) of $f_{1}$ for $v_{0}$ does not block $M^{\prime}$, we get that $u_{1}$ is saturated in $M^{\prime}$ with copies that have level at least 00. Hence, $e_{1}$ has level at most 00 for $v_{1}$ in any copy that has positive value in $M^{\prime}$. Also, if the level 00 copy of $e_{1}$ has positive value in $M^{\prime}$, then we must have that $e_{1}\succ_{u_{1}}f_{1}$, so $\mathrm{vote}_{u_{1}}(e_{1},f_{1})=+1$.
@@
5. Suppose $e_{0}\neq\emptyset$ and $e_{0}\neq e_{j_{l}}$. This implies that $\phi_{u_{0}}(e_{0},\emptyset)>0$. By the fourth axiom on sensible pairings, this implies that $\sum_{e\in E(u_{0})}N(e)<1$, so $u_{0}\notin C$, as both $M$ and $N$ saturate critical vertices. Hence, $u_{0}$ has only copies with level at least 00 in $G^{\prime}$, so $e_{0}$ has only copies with level at most 00 for $v_{0}$ in $G^{\prime}$ and thus also in $M^{\prime}$.
@@
6. Suppose that $e_{j_{l}}=\emptyset$. Then, $u_{j_{l}}$ is unsaturated by $M^{\prime}$ by the third axiom on feasible pairings. By the fact that the level 00 copy (which must exist) of $f_{j_{l}}$ for $u_{j_{l}}$ does not block $M^{\prime}$, we get that $v_{j_{l}-1}$ is saturated in $M^{\prime}$ with copies that have level at least 00. Hence, $e_{j_{l}-1}$ has level at least 00 for $v_{j_{l}-1}$ in any copy that has positive value in $M^{\prime}$. Also, if the level 00 copy of $e_{j_{l}-1}$ has positive value in $M^{\prime}$, then we must have that $e_{j_{l}-1}\succ_{v_{j_{l}-1}}f_{j_{l}}$, so $\mathrm{vote}_{v_{j_{l}-1}}(e_{j_{l}-1},f_{j_{l}})=+1$.
@@
7. Similarly as in case 5, we get that $v_{j_{l}}\notin\mathcal{C}$. Therefore, any copy of $e_{j_{l}}$ in $G^{\prime}$ have level at least 00 for $v_{j_{l}}$, thus this is also true for the ones with positive value in $M^{\prime}$. $\blacksquare$
@@
From Claim 6 it follows that the occurrences of case $2$ with $q=p+1$ are at least as many as the occurrences with $q<p$. Furthermore, we also have that $\mathrm{vote}_{v_{0}}(e_{0},f_{1})+\mathrm{vote}_{u_{1}}(e_{1},f_{1})\geq 0$ and $\mathrm{vote}_{u_{j_{l}}}(e_{j_{l}},f_{j_{l}})+\mathrm{vote}_{v_{j_{l}-1}}(e_{j_{l}-1},f_{j_{l}})\geq 0$, if $e_{0}=\emptyset$ or $e_{j_{l}}=\emptyset$ by cases 4 and 6, so we can conclude that $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))\geq 0$. $\blacksquare$
@@
Therefore, we conclude that $M$ is a popular maximum weight fractional-matching. ∎
@@
Pop-MaxW-SRI is solvable in polynomial time.
@@ 4 Conclusions
In this paper we have shown that many algorithms that rely on leveling the agents, or duplicating the edges can be extended even to roommates instances. This opens possibilities to utilize this technique even further, so it would be interesting to see other frameworks, where it helps to provide desirable solutions.
5 Acknowledgements
The author thanks Tamás Király, Kavitha Telikepalli and Yu Yokoi for fruitful discussions about popular fractional-matchings. The project was supported by the Hungarian Scientific Research Fund, OTKA, Grant No. K143858, by the Momentum Grant of the Hungarian Academy of Sciences, grant number 2021-2/2021 and by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation fund, financed under the KDP-2023 funding scheme (grant number C2258525).

### v3

@@ 3.3 Popular Maximum Weight Fractional-Matching
[…]
@@
[…]
@@
Let $E^{\prime}:=\{e\in E\mid y_{u}^{*}+y_{v}^{*}=\omega(e)\}$ and $\mathcal{C}:=\{v\in V\mid y_{v}^{*}>0\}$. Then, we have that the maximum weight fractional-matchings (so each possible candidate for a popular maximum weight fractional-matching) are exactly the ones that have positive value only on edges from $E^{\prime}$ and saturate every $v\in\mathcal{C}$. Hence, we can restrict the set of edges of $G$ to $E^{\prime}$, so from now on let $E:=E^{\prime}$.
This way, we reduced the problem to finding a popular critical fractional-matching , which to recall is a fractional-matching $M$ that saturates every $v\in\mathcal{C}$ and $\Delta(M,N)\geq 0$ for any fractional-matching $N$ that saturates every vertex $v\in\mathcal{C}$.
Therefore, now it is enough to solve Crit-fPRI.
The fine-tuned first step.
We describe the fine-tuned first step of the meta-algorithm. Let $|\mathcal{C}|=s$. We keep each edge of $E$. These will be called the middle copies. If $e=(v_{i},v_{j})$ such that $v_{i},v_{j}\notin\mathcal{C}$, we do not create additional parallel edges. If $v_{i}\in\mathcal{C},v_{j}\notin\mathcal{C}$, we add copies $e^{1},\dots,e^{s}$. These are considered worst copies for $v_{i}$ in the opposite order (i.e. $e^{s}$ is the absolute worst and $e^{1}$ is best among them) and best copies for $v_{j}$ in the opposite order (i.e. $e^{s}$ is the absolute best and $e^{1}$ is the worst among them). If $v_{i}\notin\mathcal{C},v_{j}\in\mathcal{C}$, we add copies $e^{-1},\dots,e^{-s}$. These are considered worst copies for $v_{j}$ in the opposite order (i.e. $e^{-s}$ is the absolute worst and $e^{-1}$ is best among them) and best copies for $v_{j}$ in the opposite order (i.e. $e^{-s}$ is the absolute best and $e^{-1}$ is the worst among them). If $v_{i},v_{j}\in\mathcal{C}$, then we do both of the above.
Then, the preferences are created such that the best copies are best, then the middle copies and then the worst copies are worst. With respect to that, each vertex ranks the edges of the same type according to their original order. For a vertex $v$, we use the notation $B_{j}^{v}(e)$ to denote the $j$-th best copy of $e$ among the best copies, and $W_{j}^{v}(e)$ to denote the $j$-th worst copy of $e$ among the worst copies (both can be either $e^{j}$ or $e^{-j}$). We also use the notation $\mathrm{lev}_{v}(B_{j}^{v}(e))=j$, $\mathrm{lev}_{v}(e)=0$ and $\mathrm{lev}_{v}(W_{j}^{v}(e))=-j$, which is called the level of the given copy of $e$ for $v$.
@@
Crit-fPRI is solvable in polynomial time. Furthermore, the algorithm outputs an extra-popular matching among the critical matchings.
@@
Then, we have that $\Delta(M,N,\phi)=\sum\limits_{l=1}^{k}w_{l}\cdot\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))$, where $x(P^{l})=\mathrm{vote}_{u_{0}}(e_{0},\emptyset)+\mathrm{vote}_{v_{j_{l}}}(e_{j_{l}},\emptyset)$ if $e_{0}\neq e_{j_{l}}$ and 0 otherwise.. Therefore, it is enough to show that for each walk $P^{l}$, $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))\geq 0$.
@@
    If $e_{i}$ has a level $p$ copy for $v_{i}$ in $M^{\prime}$ with positive value and $e_{i+1}$ has a level $q$ copy for $v_{i+1}$ with positive value in $M^{\prime}$, then $q\leq p+1$.
@@
    If $e_{0}=\emptyset$, then each copy of $e_{1}$ with positive value in $M^{\prime}$ has level at most 0 for $v_{1}$ and if there is one with level exactly 0, then $\mathrm{vote}_{u_{1}}(e_{1},f_{1})=+1$.
@@
    If $e_{0}\neq\emptyset$ and $e_{0}\neq e_{j_{l}}$, then each copy of $e_{0}$ with positive value in $M^{\prime}$ has level at least 0 for $v_{0}$.
@@
    If $e_{j_{l}}=\emptyset$, then each copy of $e_{j_{l}-1}$ with positive value in $M^{\prime}$ has level at least 0 for $v_{j_{l}-1}$ and if there is one with level exactly 0, then $\mathrm{vote}_{v_{j_{l}-1}}(e_{j_{l}-1},f_{j_{l}})=+1$.
@@
    If $e_{j_{l}}\neq\emptyset$ and $e_{j_{l}}\neq e_{0}$, then each copy of $e_{j_{l}}$ with positive value in $M^{\prime}$ has level at least 0 for $v_{j_{l}}$.
@@
1. Suppose that $q>p+1$. Then, the copy of $f_{i+1}$ that has level $p+1$ for $v_{i}$ blocks $M^{\prime}$, because this copy is level (p+1)>q-(p+1)>-q for $u_{i+1}$, so both of them strictly prefer it to some copies of $e_{i}$ and $e_{i+1}$ with positive value in $M^{\prime}$, contradiction.
@@
2. In this case, $u_{i+1}$ has a level (p+1)-(p+1) copy of $f_{i+1}$ with positive value in $M^{\prime}$. Hence, if either $v_{i}$ or $u_{i+1}$ prefer $f_{i+1}$ to $e_{i}$ or $e_{i+1}$ respectively, then either the level $p$ or the level $p+1$ copy of $f_{i+1}$ for $v_{i}$ blocks $M^{\prime}$, contradiction.
@@
3. Suppose both $v_{i}$ and $u_{i+1}$ prefer $f_{i+1}$ to $e_{i}$ and $e_{i+1}$ respectively. Then, by the assumption, the level $p$ copy of $f_{i+1}$ for $v_{i}$ blocks $M^{\prime}$, contradiction.
@@
4. Suppose $e_{0}=\emptyset$. Then, $\phi(f_{1},\emptyset)>0$, so $v_{0}$ is unsaturated in $M^{\prime}$ by axiom 3 on sensible pairings. By the fact that the level 0 copy (which must exist) of $f_{1}$ for $v_{0}$ does not block $M^{\prime}$, we get that $u_{1}$ is saturated in $M^{\prime}$ with copies that have level at least 0. Hence, $e_{1}$ has level at most 0 for $v_{1}$ in any copy that has positive value in $M^{\prime}$. Also, if the level 0 copy of $e_{1}$ has positive value in $M^{\prime}$, then we must have that $e_{1}\succ_{u_{1}}f_{1}$, so $\mathrm{vote}_{u_{1}}(e_{1},f_{1})=+1$.
@@
5. Suppose $e_{0}\neq\emptyset$ and $e_{0}\neq e_{j_{l}}$. This implies that $\phi_{u_{0}}(e_{0},\emptyset)>0$. By axiom 4 on sensible pairings, this implies that $\sum_{e\in E(u_{0})}N(e)<1$, so $u_{0}\notin\mathcal{C}$, as both $M$ and $N$ saturate critical vertices. Hence, $u_{0}$ has only copies with level at least 0 in $G^{\prime}$, so $e_{0}$ has only copies with level at most 0 for $v_{0}$ in $G^{\prime}$ and thus also in $M^{\prime}$.
@@
6. Suppose that $e_{j_{l}}=\emptyset$. Then, $u_{j_{l}}$ is unsaturated by $M^{\prime}$ by axiom 3 of feasible pairings. By the fact that the level 0 copy (which must exist) of $f_{j_{l}}$ for $u_{j_{l}}$ does not block $M^{\prime}$, we get that $v_{j_{l}-1}$ is saturated in $M^{\prime}$ with copies that have level at least 0. Hence, $e_{j_{l}-1}$ has level at least 0 for $v_{j_{l}-1}$ in any copy that has positive value in $M^{\prime}$. Also, if the level 0 copy of $e_{j_{l}-1}$ has positive value in $M^{\prime}$, then we must have that $e_{j_{l}-1}\succ_{v_{j_{l}-1}}f_{j_{l}}$, so $\mathrm{vote}_{v_{j_{l}-1}}(e_{j_{l}-1},f_{j_{l}})=+1$.
@@
7. Similarly as in case 5, we get that $v_{j_{l}}\notin\mathcal{C}$. Therefore, any copy of $e_{j_{l}}$ in $G^{\prime}$ have level at least 0 for $v_{j_{l}}$, thus this is also true for the ones with positive value in $M^{\prime}$.  $\blacksquare$
@@
From Claim 9 cases 4 -7 it follows that the occurrences of case 2 with $q=p+1$ are at least as many as the occurrences with $q<p$. Furthermore, we also have that $\mathrm{vote}_{v_{0}}(e_{0},f_{1})+\mathrm{vote}_{u_{1}}(e_{1},f_{1})\geq 0$ and $\mathrm{vote}_{u_{j_{l}}}(e_{j_{l}},f_{j_{l}})+\mathrm{vote}_{v_{j_{l}-1}}(e_{j_{l}-1},f_{j_{l}})\geq 0$, if $e_{0}=\emptyset$ or $e_{j_{l}}=\emptyset$ by cases 4 and 6, so we can conclude that $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))\geq 0$ (by case 2 and $\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})\geq-2$ even if $q<p$).  $\blacksquare$
@@
Therefore, we have that $M$ is a popular maximum weight fractional-matching. ∎
@@
MaxW-fPRI is solvable in polynomial time.
@@ 4 Conclusions
In this paper we have shown that many algorithms that rely on leveling the agents, or duplicating the edges can be extended even to roommates instances. We extended the $3/2$-approximation algorithm for maximum-size weakly stable matchings from bipartite graphs to arbitrary instances. This resolved a long-standing open question and has important applications, particularly in allocating junior doctors to hospitals in the presence of couples among the doctors.
Our central observation may open even more possibilities to utilize this technique even further, so it would be interesting to see other frameworks, where it helps to provide desirable solutions.